### v1

# Ultrafast optical excitation of magnons in 2D antiferromagnets via spin torque exerted by photocurrent of excitons: Signatures in charge pumping and THz emission

Abstract
Recent experiments observing femtosecond laser pulse (fsLP) exciting magnons within two-dimensional (2D) antiferromagnetic (AF) semiconductors—such as CrSBr, NiPS3, and MnPS3, or their van der Waals heterostructures—suggest exciton-mediation of such an effect. However, its microscopic details remain obscure as resonant coupling of magnons, living in the sub-meV energy range, to excitons, living in $\sim 1$ eV range, can hardly be operative. Here, we develop a quantum transport theory of this effect, in which time-dependent nonequilibrium Green’s function (TDNEF) for electrons driven by classical vector potential of fsLP are coupled to the Landau-Lifshitz-Gilbert (LLG) equation describing classical dynamics of localized magnetic moments (LMMs) within 2D AF semiconductor. Our TDNEGF+LLG theory explains how fsLP, with central frequency above the semiconductor gap, generates a photocurrent that subsequently exerts spin-transfer torque (STT) onto LMMs as a genuine nonequilibrium spintronic effect. The collective motion of LMMs analyzed by windowed Fast Fourier transform (FFT) reveals frequencies of excited magnons, as well as their lifetime governed by nonlocal damping due to the bath of electrons. In addition, the TDNEGF part of our TDNEGF+LLG self-consistent loop computes a time-dependent density matrix whose off-diagonal elements are utilized to describe, at the mean-field level, inter-orbital Coulomb interaction binding electrons and holes into excitons. Our TDNEGF+LLG theory predicts how excited magnons pump charge current into the attached electrodes, or locally within AF semiconductor responsible for microwave emission. The windowed FFT of the former signal contains imprints of excited magnons, as well as their interaction with excitons, which could be exploited as a novel probe in future experiments.
Introduction.—The advent of two-dimensional (2D) magnetic materials [1, 2]—such as 2D antiferromagnetic (AF) semiconductors CrSBr [3, 4, 5, 6, 7, 8, 9], NiPS3 [10], MnPS3 [11] and their van der Waals heterostructures [12, 13]—has made possible recent experiments observing how femtosecond laser pulse (fsLP) excites magnons with presumed exciton mediatation. From a fundamental viewpoint, these experiments rekindle [14] interest in exciton-magnon coupling that was dormant  [^1] Studies [16, 17, 18, 19, 20, 21, 22] of exciton-magnon coupled systems in the 1960s and 1970s were largely motivated by absorption spectra of particular three-dimensional AF insulators like MnF2. for many decades [16, 17, 18, 19, 20, 21, 22]. Such coupling can lead to intriguing quantum many-body effects, including magnon-magnon interactions [23, 24] dressed by excitons [5], or magnon-mediated exciton-exciton interactions [9, 25]. As regards applications, in magnonics [26] for classical information processing, there is a considerable effort to excite coherent magnons [27, 28] by ultrafast light, whose frequencies are as high as possible and wavelengths as short as possible [29, 11]. This is because magnons with wavelength $\lesssim 100$ nm would enable miniaturization of envisaged magnonic devices down to the nanoscale [30, 31]. In contrast, oscillating magnetic fields (supplied via microstrip lines or coplanar waveguides) as standard tools are impractical for AF materials. Other schemes demonstrating excitation of AF magnons, such as by injecting current [32], lead to diffusive propagation of incoherent (at many uncontrolled frequencies) magnons. Furthermore, for quantum information processing exciton-magnon coupling offers potential of transduction [33, 34, 2] of quantum information from qubits to microwaves that would excite magnons, and then from them, via excitons, further transduction to optical photons. In turn, photons can transfer quantum information over long distances via optical fibers [34].
Figure 1: Schematic view of a two-terminal setup for TDNEGF+LLG [Fig. 2] calculations of nonequilibirum dynamics of photoexcited electrons coupled to local magnetization. The central active region (CAR) consists of two layers of 2D AF semicoductor which are described by the top and bottom TB chains (as inspired by quasi-1D structure of CrSBr [35, 36]) hosting two orbitals, $c$(onduction) and $v$(alence), per site $i$. Electron spin densities (green arrows), $\langle\hat{\mathbf{s}}_{i_{c}}(t)\rangle$ and $\langle\hat{\mathbf{s}}_{i_{v}}(t)\rangle$, interact via $sd$ exchange [Eq. (10) and (11)] with classical LMM at the same site described by vector (red arrow) $\mathbf{M}_{i}(t)$ obeying the LLG Eq. (3). Electrons on $c$ and $v$ orbitals at the site $i$ interact via inter-orbital local Coulomb interaction of strength $U$ [Eq. (14)] which, when turned on $U>0$, binds photoexcited electrons and holes into excitons [37, 38]. Both chains are attached to semi-infinite ideal NM leads modeled as 1D TB chains. Via such leads, any spin or charge current pumped within CAR by fsLP, or by excited magnons [39, 40, 41] persisting after fsLP ceases, is drained [Fig. 4] toward macroscopic reservoirs kept at the same Fermi energy $E_{F}$ (i.e., no bias voltage is applied between the leads). We also compute EM radiation, emitted by pumped local charge currents within the CAR, and analyze its frequency content in the THz range [Fig. 5].
The possibility of precise tuning of the central frequency of fsLP around subgap electronic states (like exciton or on-site $d$-$d$ transition [42, 43]) and thereby achieved control of excited magnons (such as dependence on light polarization [43]) emphasizes the key role [44] played by photoexcited electrons as mediators of magnon excitation. Thus, electrons capable of responding fast to fsLP must be explicitly included in any microscopic theory [45]. Their inclusion displaces often invoked [46, 47, 48, 49], via intuitive reasoning, direct coupling of light to local magnetization of gapped materials which is typically negligible effect when examined via first-principles calculations [50]. However, the first-principles derived Hamiltonians of exciton-magnon interactions are currently lacking [51]. Furthermore, extraction of magnon and exciton spectra from a single first-principles framework (as possible via GW methodology [52]) for a chosen 2D magnetic material often shows their vastly different energy scales [52], meaning that any exciton-magnon coupling cannot be resonant.
Lacking such an input Hamiltonian, the magnon excitation aspect of recent experiments is typically accounted for phenomenologically [3, 4, 5, 13, 8] via the Landau-Lifshitz-Gilbert (LLG) equation [53] for localized magnetic moments (LMMs) viewed as classical (unit) vectors $\mathbf{M}_{i}(t)$ [Fig. 1]. While no effect of photoexcited electrons is explicitly included in the LLG equation, it has been conjectured that configuration of LMMs can affect exciton energy in the case of CrSBr [3, 4, 5]. The agreement between such phenomenological models and experimentally excited magnon spectra implies that magnons can be approximately treated as classical spin waves [54, 55], despite challenges [56, 57] that AF materials pose for classical LLG treatment. We recall that in LLG description, magnons [54, 55, 24] emerge as collective excitations above a magnetically ordered ground state in which $\mathbf{M}_{i}(t)$ precess around a direction specified by magnetic anisotropy and/or externally applied magnetic field [3, 4, 6, 7, 8, 5]. The phase of precession $\mathbf{M}_{i}(t)$ of adjacent vectors varies harmonically in space over the magnon wavelength $\lambda$. However, microscopic understanding is lacking regarding what provides the “kick” in the LLG equation (e.g., Refs. [5, 13] assume presence of some type of phenomenological “impulsive perturbation”). Furthermore, excited magnons could imprint their signatures [40, 41, 39] on currents flowing through 2D AF semiconductor, but their properties and their usage as novel experimental probe remain unexplored.
Figure 2: Flowchart of TDNEGF+LLG self-consistent loop combining TDNEGF [58, 59, 60] (green box) computation [Eq. (4)] of time-dependent nonequilibrium density matrix ${\bm{\rho}}^{\rm neq}(t)$ with LLG equation [53, 55, 54] updating (orange box) LMMs $\mathbf{M}_{i}(t)$. For this study, to the previously developed [61, 62] TDNEGF+LLG scheme, we add a computation (magenta box) employing [38] all off-diagonal elements of ${\bm{\rho}}^{\rm neq}(t)$ in order to describe the binding of photoexcited conduction electrons and valence holes into exctions. The loop employs time step $\delta t=0.1$ fs in both quantum (as required for numerical stability of TDNEGF calculations [58, 59, 61]) and classical LLG calculations. After each time step, we obtain time-dependent observables [Eqs. (1), (6), (7) and (17)] listed in the yellow box. In particular, STT is constructed from the expectation value of electron spin $\langle\hat{\mathbf{s}}_{i}(t)\rangle$ [Eq. (1)] and $\mathbf{M}_{i}(t)$ via Eq. (2).
In this Letter, we also employ the LLG equation, but we introduce the “kick” microscopically by self-consistently coupling LLG equation to computational quantum transport [63] of photoexcited electrons. Such electrons are described by the time-dependent nonequilibrium Green’s function (TDNEGF) formalism [58, 59, 60], as illustrated in Fig. 2. Thus, TDNEGF+LLG framework [61, 62, 39, 64, 65] makes it possible to introduce into the LLG equation, in numerically exact fashion, the effect of photocurrent [66] ignited by fsLP. The photocurrent will become spin-polarized as it propagates through the magnetic environment created by LMMs. In addition, since LMMs within AF CAR in Fig. 1 will be noncollinear due to inevitable thermal fluctuations  [67, 68] or applied magnetic field in experiments [3, 4, 5] that we also include [Eq. (10)], the nonequilibrium spin density
$\langle\hat{\mathbf{s}}_{ia}\rangle(t)=\mathrm{Tr}_{\mathrm{spin}}[{\bm{\rho}}^{\rm neq}(t){\bm{\sigma}}],$ (1)
of photoexcited electrons will lead to nonzero spin-transfer torque (STT) [69, 70, 71] on the LMMs $\mathbf{M}_{i}$,
$\mathbf{T}_{i}(t)=J_{sd}\left[\sum_{a=c,v}\langle\hat{\mathbf{s}}_{ia}\rangle(t)\right]\times\mathbf{M}_{i}(t).$ (2)
The STT describes [69, 70, 71] how flowing electrons transfer spin angular momentum to local magnetization through $sd$ exchange interaction $J_{sd}$ [72]. Here $a=c,v$ labels each of two orbitals per site [Fig. 1]; ${\bm{\sigma}}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ is the vector of the Pauli matrices; ${\bm{\rho}}^{\rm neq}(t)$ is time-dependent nonequilibrium density matrix [58, 59, 73]; and trace $\mathrm{Tr}_{\mathrm{spin}}[\ldots]$ is over quantum states in the spin space only. Such microscopically computed STT is sent into the LLG equation,
$\partial_{t}\mathbf{M}_{i}$ $=$ $-g_{0}\mathbf{M}_{i}\times\mathbf{B}^{\mathrm{eff}}_{i}+\alpha_{G}\mathbf{M}_{i}\times\partial_{t}\mathbf{M}_{i}+\frac{g_{0}}{\mu_{M}}\mathbf{T}_{i},$ (3)
while dynamics of $\mathbf{M}_{i}(t)$ modifies the quantum Hamiltonian [Eq. (11)] of electrons within the loop in Fig. 2. This then establishes self-consistency between TDNEGF and LLG calculations, as initially photoexcited spin current, $I^{S_{\alpha}}=\mathrm{Tr}[{\bm{\rho}}^{\rm neq}(t)\hat{I}^{S_{\alpha}}]$, will be modified by the dynamics of $\mathbf{M}_{i}(t)$; whereas their trajectories are, in turn, affected by updated spin current and STT $\mathbf{T}_{i}(t)$ exerted by it. In LLG Eq. (3) $g_{0}$ is gyromagnetic factor [53]; $\mu_{M}$ is the magnitude of LMMs [53]; $\mathbf{B}^{\rm eff}_{i}=-\frac{1}{\mu_{M}}\partial\mathcal{H}/\partial\mathbf{M}_{i}$ is effective magnetic field obtained from classical Hamiltonian $\mathcal{H}$ for LMMs [36, 74]; Gilbert damping is chosen as $\alpha_{G}=0.01$ as the typical value for Cr-based 2D magnets [74]; and we use shorthand notation $\partial_{t}\equiv\partial/\partial t$.
Figure 3: Time dependence of the Néel vector, defined [Eq. (8)] for two monolayers of AF semiconductor CAR [Fig. 1] in: (a) the absence of excitons ($U=0$); or (c) their presence induced by on-site Coulomb interaction [37, 38] $U=0.1\gamma$ [Eq. (14)]. Panels (b) and (d) plot the corresponding power spectrum of windowed FFT [75, 76], $|N^{y}(t_{c},\omega)|^{2}$, revealing frequencies and lifetime of magnons excited by STT from photocurrent of conduction electrons in (a) or excitons in (b), respectively. Here $t_{c}$ denotes the central time of the Gaussian window [75, 76] used in FFT. Note that $\omega_{b}$ labels the frequency of the bright magnon, which corresponds to the same type of magnon in bilayer AF semiconductors observed in experiments of Refs. [3, 4, 5]. Gray curves on the bottom of panels (a) and (c) depict vector potential $A^{x}(t)$ of fsLP.
The fundamental quantity of quantum statistical mechanics is the density matrix. The time-dependent one-particle nonequilibrium density matrix, ${\bm{\rho}^{\rm neq}}(t)=\hbar\mathbf{G}^{<}(t,t)/i$, can be expressed in terms of the lesser Green’s function of TDNEGF formalism $G^{<,\sigma\sigma^{\prime}}_{ii^{\prime}}(t,t^{\prime})=\frac{i}{\hbar}\langle\hat{c}^{\dagger}_{i^{\prime}\sigma^{\prime}}(t^{\prime})\hat{c}_{i\sigma}(t)\rangle_{\mathrm{nes}}$  [58] where $\langle\ldots\rangle_{\mathrm{nes}}$ is the nonequilibrium statistical average [60]. We solve a matrix integro-differential equation [59]
$i\hbar\partial_{t}{\bm{\rho}}^{\rm neq}=[\mathbf{H}(t),{\bm{\rho}}^{\rm neq}]+i\sum_{p=L,R}[{\bm{\Pi}}_{p}(t)+{\bm{\Pi}}_{p}^{\dagger}(t)],$ (4)
for the time evolution of ${\bm{\rho}}^{\rm neq}(t)$, where $\mathbf{H}(t)$ is the matrix representation of the quantum Hamiltonian of electrons. Equation (4) is an exact quantum master equation for the reduced density matrix of the AF CAR in Fig. 2 viewed as an open finite-size quantum system attached to macroscopic Fermi liquid reservoirs via semi-infinite normal metal (NM) leads. The NM leads, not used in experiments [3, 4, 6, 7, 8, 5, 9] thus far, are important technically within TDNEGF calculations in order to introduce continuous energy spectrum and dissipation effects  [77], thereby guaranteeing that excited photocurrent and $\mathbf{M}_{i}(t)$ dynamics will eventually cease at some time after fsLP. Otherwise, in a closed quantum system [68] without surrounding bath [67] one would find forever oscillating photocurrent, which is obviously unphysical. Furthermore, the leads allow us to analyze properties of charge and spin currents outflowing into them. Such currents could also offer novel experimental probe of excitons, magnons and their interactions, as we confirm in Fig. 5. For this purpose, we use ${\bm{\Pi}}_{p}(t)$ matrices
[…]
expressed in terms of the lesser and greater Green’s functions [60] and the corresponding self-energies ${\bm{\Sigma}}_{p}^{>,<}(t,t^{\prime})$ [59], to obtain time-dependent charge
$I_{p}(t)=\frac{e}{\hbar}\mathrm{Tr}\,[{\bm{\Pi}}_{p}(t)],$ (6)
and spin
$I_{p}^{S_{\alpha}}(t)=\frac{e}{\hbar}\mathrm{Tr}\,[\hat{\sigma}_{\alpha}{\bm{\Pi}}_{p}(t)],$ (7)
currents outflowing into $p=L,R$ NM leads. Since the applied bias voltage between the left ($L$) and right ($R$) NM leads is identically zero in the setup of Fig. 1, all computed currents $I_{p}(t)$ and $I_{p}^{S_{\alpha}}(t)$ are solely due to pumping [78, 79, 61, 64, 80, 39] by (nonperiodic [78]) time-dependence of the CAR Hamiltonian. Note that we use the same units for charge and spin currents, defined as $I_{p}=I_{p}^{\uparrow}+I_{p}^{\downarrow}$ and $I_{p}^{S_{\alpha}}=I_{p}^{\uparrow}-I_{p}^{\downarrow}$, in terms of spin-resolved charge currents $I_{p}^{\sigma}$. In our convention, positive current in NM lead $p$ means charge or spin current is flowing out of that NM lead.
Figure 4: Time-dependence of charge current $I_{R}(t)$ flowing into the right NM lead in Fig. 1 due to pumping  [39, 41, 40] by excited magnons from Fig. 3 in: (a) the absence of excitons ($U=0$); or (c) their presence induced by on-site Coulomb interaction $U=0.1\gamma$ [Eq. (14)]. Panels (b) and (d) plot the corresponding power spectrum of windowed FFT [75, 76], $|I_{R}(t_{c},\omega)|^{2}$. Gray curves on the bottom of panels (a) and (c) depict vector potential $A^{x}(t)$ of fsLP.
Let us recall that problem of how STT excites uniform motion of all LMMs [i.e., of their macrospin $\mathbf{M}(t)=\sum_{i}\mathbf{M}_{i}(t)$] vs. their nonuniform motion like magnons was analyzed long ago [81] for conventional metallic ferromagnets, as well as recently for AF insulators [82, 65], with focus on threshold injected current value [32, 83] for magnons to occur. Our setup in Fig. 2 is different from those studies, as no current is injected from an external circuit. Instead, photocurrent is excited within the AF semiconductor CAR by fsLP, and can eventually exit into the NM leads. In addition to photocurrent of conduction electrons, we also consider photocurrent of excitons generated by Coulomb interaction, which binds together electrons in the conduction band with holes in the valence band. To capture such binding, we employ a time-dependent mean-field theory (tMFT) [38, 37] of inter-orbital Coulomb interaction term [Eq. (14)]. For this purpose, we exploit [38] the off-diagonal elements of $\mathbf{\bm{\rho}}^{\rm neq}(t)$, that we naturally construct within TDNEGF part [59, 58] of TDNEGF+LLG self-consistent loop [61, 62, 39, 64, 65]. In other words, for the study presented here, the previously developed [61, 62, 39, 64, 65, 84] TDNEGF+LLG framework is amended with modeling of excitonic effects. Details of our quantum Hamiltonian for the electronic subsystem, $\hat{H}$, and classical one, $\mathcal{H}$, for the subsystem of LMMs, together with explanation of tMFT and electromagnetic radiation (EM) calculations via the Jefimenko equations [85, 86, 87, 88, 89, 78, 90], are provided in End Matter.
Figure 5: Time-dependence of $E_{\mathrm{FF}}^{x}(t)$ the component of the electric field in FF region of EM radiation emitted by bond charge currents [Eq. (17)] within the CAR of Fig. 1 in: (a) the absence of excitons ($U=0$); or (c) their presence induced by on-site Coulomb interaction $U=0.1\gamma$ [Eq. (14)]. Panels (b) and (d) plot the corresponding power spectrum of windowed FFT [75, 76], $|E_{\mathrm{FF}}^{x}(t_{c},\omega)|^{2}$. Gray curves on the bottom of panels (a) and (c) depict vector potential $A^{x}(t)$ of fsLP.
Results and Discussion.—We first recall definition [3, 4, 5, 8] of the Néel vector
$\mathbf{N}\equiv(N^{x},N^{y},N^{z})=\frac{1}{2N}\sum_{i\in 1,j\in 2}\big{(}\mathbf{M}_{i}-\mathbf{M}_{j}\big{)},$ (8)
between two monolayers of AF semiconductor, where in equilibrium $\mathbf{N}(t=0)\equiv(2,0,0)$. Out of equilibrium, as initiated by fsLP, the Néel vector starts evolving in time [Figs. 3(a) and 3(c)]. Somewhat surprisingly and also observed experimentally [13], this evolution starts without substantial delay with respect to fsLP, even though LMMs are slower than electrons [67]. Because the magnon frequency spectrum encoded in $N^{z}(t)$ could be changing within different time frames, we apply windowed (or short-time) Fast Fourier transform (FFT) [75, 76, 91, 92]. For example, magnons get excited around $t_{c}\simeq 15\ \mathrm{ps}$ in Fig. 3(b), and subsequently they decay [23, 26] because of Gilbert damping $\alpha_{G}$ in Eq. (3). Furthermore, because of explicitly introduced electrons via TDNEGF calculations, additional nonlocal damping  [62, 64, 93] is introduced into Eq. (3) via the STT $\mathbf{T}_{i}(t)$ term. Due to such decay, excited magnons vanish at around $t_{c}\simeq 18\ \mathrm{ps}$ in Fig. 3(b). For such signals—appearing also in many other scientific disciplines (such as electroencephalography [76] or speech analysis)—it is advantageous to perform windowed FFT over successive time intervals. For this purpose, we employ Gaussian as the window function,
[…]
which makes it possible to extract time-frequency content from signals whose oscillations are localized in a finite time frame. Here, $t_{c}$ is the centroid of the Gaussian, serving as the abscissa of panels (b) and (d) within each of Figs. 35 analyzing $X(t)=N^{\alpha}(t),I_{R}(t),E^{x}_{\mathrm{FF}}(t)$ as the signal, respectively. Here, $\sigma$ specifies the width of the Gaussian and $\lambda$ is a parameter controlling the resolution. For example, greater values of $\lambda\sigma$ yield better resolution in the time domain, while smaller values yield improved resolution in the frequency domain, where the frequency and time resolution satisfy a Heisenberg-like uncertainty relation [75, 76]. Windowed FFT of the Néel vector produces $N^{z}(\omega,t_{c})$, whose power spectrum in Figs. 3(b) and 3(d) reveals excitation of the so-called “bright magnon” at the frequency $\omega_{b}$. Thus, our TDNEGF+LLG theory explains fully microscopically the same “bright magnon” observed experimentally [3, 4, 5, 8]. In the presence of excitons due to nonzero $U$, we find longer lifetime of excited magnons in Fig. 3(d). Both Figs. 3(b) and  3(d) also show short (for $U=0$) vs. longer (for $U\neq 0$) living magnons, respectively, that are excited at frequencies below $\omega_{b}$. We confirm in Fig. S1(a) of the Supplemental Material (SM) 222 See Supplemental Material at https://wiki.physics.udel.edu/qttg/Publications for an additional Fig. S1(a) that is the counterpart of Fig. 3(b) for the case when photocurrent is allowed to excite magnons via STT, but then $sd$ exchange interaction between electron spin and LMMs is quickly switched off. For easy comparison, we also analyze in Fig. S1(b) experimental data from Ref. [5] using windowed FFT. that such magnons are a direct consequence of $J_{sd}\neq 0$.
The excited magnons will introduce the second nonequlibrium drive into the subsystem of electrons. Such time-dependent drive can lead to pumping [40, 39, 41] of electronic spin and charge currents. They can be differentiated from currents pumped [66] by fsLP, as the first nonequilibirum drive, by their specific frequency content. We compute charge current $I_{R}(t)$ [Figs. 4(a) and 4(c)] pumped in the right NM lead, and perform windowed FFT on it to obtain $|I_{R}(t_{c},\omega)|^{2}$ and plot it [Figs. 4(b) and 4(d)] in the same frequency range where magnons are found in Fig. 3. This power spectrum contains peaks at $\omega_{b}$, meaning that by attaching an additional external circuit to recent experiments and by analyzing pumped current into that circuit, one could confirm optical excitations of magnons. Note that pumped current is also a sensitive probe for the formation of excitons, as in $U\neq 0$ case long-lived high harmonic generation (HHG) emerges in Fig. 4(d).
Finally, we examine EM radiation that will be emitted by time-dependent local (or bond [95]) charge currents $I_{ia\rightarrow jb}(t)$ [Eq. (17)] within AF semiconductor CAR in Fig. 1. Note that EM radiation emitted by ultrafast-light-driven magnetic materials and their heterostructures is routinely used [96, 97, 98, 99] in spintronic experiments as a probe of coupled spin-charge dynamics [89, 100, 78, 90] of such far-from-equilibrium systems. We compute the $x$-component $E^{x}_{\mathrm{FF}}(\mathbf{r},t)$ of electric field of EM radiation in the far-field (FF) region—where radiation decays as $\sim 1/r$, such as by using $\mathbf{r}_{0}=(5a,0,1000a)$—via the Jefimenko [85, 87] Eq. (16) using $I_{ia\rightarrow jb}(t)$ as the source [88, 67]. The windowed FFT of $E^{x}_{\mathrm{FF}}(\mathbf{r}_{0},t)$ in Figs. 5(a) and 5(c) yields $E^{x}_{\mathrm{FF}}(\mathbf{r}_{0},t_{c},\omega)$ whose power spectrum is plotted in Figs. 5(b) and 5(d), respectively. Similarly to $|I_{R}(t_{c},\omega)|^{2}$, the power spectrum $|E^{x}_{\rm FF}(t_{c},\omega)|^{2}$ contains an imprint of excited “bright magnon”, as well as HHG signaling [Fig. 5(d)] the presence of exciton-magnon interactions. Note that the frequency content of EM radiation related to magnons is in the range of: THz in Fig. 5 for our model; sub-THz for CrI3 in Ref. [13]; and GHz for CrSBr in Ref. [5].
Conclusions and Outlook.—Using computational time-dependent quantum transport [58, 59], extended [38, 37] to include binding of photoexcited electrons and holes into excitons, we demonstrate that microscopic mechanism behind recent experimental observations [3, 4, 5, 13, 43, 8] of optical excitation of “bright magnon” in bilayer 2D AF semiconductors is a nonequilibrium spintronic effect of spin torque type [69, 101]. This approach displaces the need for phenomenological picture of “impulsive perturbation” [5, 13], invoked within the classical LLG equation alone to interpret these experiments, while revealing additional signatures of magnons and their interaction with excitons that could be exploited in future experiments. We note that very recent experiments [5] on CrSBr have also observed HHG of magnons at frequencies $n\omega_{b}$, with $n$ reaching surprisingly large $n\gtrsim 20$. In contrast, our theory [Fig. 3] finds only “bright magnon” at frequency $\omega_{b}$. Although the LLG equation is nonlinear and can, in principle, capture [57, 24] magnon-magnon interactions that are likely one of the key ingredients [102] for large $n$, it seems that LLG dynamics on its own is insufficient. We anticipate that adding quantum corrections [103, 104] to LLG equation, as often required for AF materials [56, 57, 102], and/or by treating excitons beyond tMFT which can be achieved [37, 60] within TDNEGF formalism by including additional self-energies into our Eq. (4), could explain HHG of magnons in experiments [5]. We relegate such extensions of TDNEGF+LLG framework and applications to this unsolved problem to future studies.
Acknowledgements.
J. V.-M. and Y. R. were supported by the University of Delaware (UD) Research Foundation Strategic Initiative Award. Y. R. and B. K. N. were supported by the US National Science Foundation (NSF) through the UD Materials Research Science and Engineering Center (MRSEC), DMR-2011824. The supercomputing time was provided by DARWIN (Delaware Advanced Research Workforce and Innovation Network), which is supported by NSF Grant No. MRI-1919839.